OpenStudy (anonymous):

whats the answer for (6x3-4x+2)-(2x3-1)

OpenStudy (anonymous):

do you want it simplified?

OpenStudy (anonymous):

and do the x's stand for a varible or multiplication?

OpenStudy (anonymous):

\[8 \times x^{3} - 4 \times x +3\]

OpenStudy (anonymous):

x=multiplication: (6x3-4x+2)-(2x3-1) (18-4x +2)-(6-1) (16-4x) - (5) 11-4x

OpenStudy (mertsj):

\[6x^3-4x+2-(2x^3-1)=6x^3-4x+2-2x^3+1=4x^3-4x+3\]

OpenStudy (mertsj):

So the answer is 4x^3-4x+3